# Electromagnetic energy calibration of the SoLid detector with horizontal muons

### v3

@@ 6.2 The linearity curve
It is interesting to gather the different calibration results obtained with the SoLid detector in a single linearity plot. The plot shows the calibration values obtained from the horizontal cosmic muons, the cosmogenic background source 12B, and the AmBe radioactive source measurements covered in this paper. It is complemented with an additional point provided by another radioactive source 22Na, which was previously used by the collaboration and is discussed extensively in [20]. A coherent picture emerges from this figure and provides confidence in the standard calibration tools of the detector response used for the physics measurements discussed in [21].

### v4

@@ 6.2 The linearity curve
It is interesting to gather the different calibration results obtained with the SoLid detector in a single linearity plot. The plot shows the calibration values obtained from the horizontal cosmic muons, the cosmogenic background source 12B, and the AmBe radioactive source measurements covered in this paper. It is complemented with an additional point provided by another radioactive source 22Na, which was previously used by the collaboration and is discussed extensively in Ref. [21]. A coherent picture emerges from this figure and provides confidence in the standard calibration tools of the detector response used for the physics measurements discussed in Ref. [22].